(1 point) Find the Laplace Transform of the following functions: f(t) = 2e-9t + 7++ 4t+3 F(s) = f(t) = 2e9t sin(7t) + 4ť + 3e1
Answer #1

Here, tat? 76 +72lt²) (778t (777ht + 3 S = + zet 2 I ( ate-gtzin (It) fet) = 2e-9t +43+3 :: F(5) F(5) = b (f(t)) =2L (e-94) s